OpenStudy (anonymous):

Factor 6(x − 4)2 − (x − 4) − 2 completely.

OpenStudy (anonymous):

=12x+−48+−x+4+−2 =12x+−48+−x+4+−2 =(12x+−x)+(−48+4+−2) =11x+−46 Answer: =11x−46

OpenStudy (anonymous):

thanks alot

OpenStudy (campbell_st):

wow... and I thought the problem was \[6(x -4)^2 - (x -4) - 2\] which is an equation reducible to a quadratic which can be factored (6(x -4) - 4)(6(x -4) + 3) --------------------- 6 and then simplifies to (3(x - 4) -2)(2(x - 4) + 1) or after ditributing inside each factor (3x -14)(2x -7) but them maybe it was the simple solution shown above